Please:
My question is to findout what is wrong with this:
(-1)^2 = (1)^2
log(-1)^2 = log(1)^2
2log(-1) = 2log(1)
(-1) = (1) ???????

Answer by lwsshak3(11628)

Please:
My question is to findout what is wrong with this:
(-1)^2 = (1)^2
log(-1)^2 = log(1)^2
2log(-1) = 2log(1)
(-1) = (1) ???????
**
You cannot take the log of a negative number:
Domain of logarithms: (0,∞), that is, the number must be>0